OpenStudy (anonymous):

((3)^-2)^x=√(3) Find x. Also, if possible, please explain how you got the answer. Thanks!

OpenStudy (anonymous):

(3^(-2))^x = 3^(1/2) 3^(-2x) = 3^(1/2) -2x = 1/2 x=-1/4 main formula used is that (x^y)^z) = x^(yz)

OpenStudy (anonymous):

Thanks so much! :D